Operations Report – February 22th

Crew 275 Operations Report 22Feb2023
SOL: 10

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9

Notes on non-nominal systems:
• See section “Summary of suits and radios” below

Spirit rover used: No
• Hours: 209.1
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: Yes
• Hours: 113.2
• Beginning charge: 100%
• Ending charge: 70%
• Currently charging: Yes

Curiosity rover used: No
• Hours: 221.0
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: Yes
• Hours: 255.6
• Beginning charge: 100%
• Ending charge: 61%
• Currently charging: Yes

General notes on rovers: Rovers were used during today’s EVA. Nothing to report.

Summary of Hab operations: The Crew conducted routine operations. Due to high winds, the main Airlock door and the Engineering Airlock doors were closed and secured.
• WATER USE: 13.90 gallons
• Water: 329.08 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS.

Summary of suits and radios: During EVA preparation, as a Crew Member was checking the battery of the Spacesuit n°9, he accidentally opened too high the back wooden panel. The panel was subsequently broken, and the panel separated from the spacesuit, breaking at the hinges. I can provide additional photos if you need.

Summary of GreenHab operations: The Crew botanist conducted routine operations.
• WATER USE: 10.57 gallons
• Heater: On
• Supplemental light: No
• Harvest: 65 g of beans – 3 g of mint – 2 g of thyme

Summary of ScienceDome operations: The Crew Scientist conducted routine operations. The preparation of the EVA concerning the photogrammetry experiment was performed.
• Dual split: Switched off during the night

Summary of RAM operations: No operations were conducted in the RAM today.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Today’s EVA was aborted after 80 minutes on field, because of the very high winds and the cold temperature felt at North Ridge.

Questions, concerns, and requests to Mission Support:
• Would it be possible to have a supplemental table and a supplemental chair in the GreenHab? Indeed, we need to perform psychometric tests there.

Copyright © The Mars Society. All rights reserved. | Main Site